# MENTOR: Efficient Multimodal-Conditioned Tuning for Autoregressive Vision Generation Models

### v1

@@ E.1 Benchmark Details
Data Organization. DreamBench++ [Peng et al., 2025] comprises 150 high-quality reference images, sourced from Unsplash, Rawpixel, and Google Images, encompassing a balanced mix of subjects. These are evenly divided into three broad categories: objects, living subjects (humans and animals), and styles (illustrative, painterly, etc.), ensuring visual and conceptual diversity.
In total, DreamBench++ offers 1,350 prompts ($150\times 9$), representing a substantial scale-up over the original DreamBench (30 subjects $\times$ 25 prompts). Relative to DreamBench, the dataset is 5$\times$ larger in subjects and 54$\times$ larger in prompts, enabling broader evaluation of generative performance.
@@
The original DreamBench [Ruiz et al., 2023] dataset consists of 30 subjects, each paired with 25 prompts, totaling 750 prompt-image pairs. It serves as a foundational benchmark for evaluating personalized image generation models, focusing on the model’s ability to maintain subject identity across diverse prompts.
@@ E.2 Baselines
  • Textual Inversion [Gal et al., 2022] learns a new word embedding to represent a specific concept, enabling personalized image generation by incorporating the new token into prompts. It requires a few images of the subject and fine-tunes the embedding without altering the base model weights.
@@
  • DreamBooth [Ruiz et al., 2023]: DreamBooth fine-tunes a pre-trained text-to-image model to bind a unique identifier with the subject’s visual concept, allowing for personalized generation. It requires several images of subject and modifies model weights to capture subject-specific details.
@@
  • BLIP-Diffusion [Li et al., 2023a]: This approach introduces a pre-trained multimodal encoder to provide subject representations for the diffusion generator, enabling controllable multimodal image generation.
@@
  • KOSMOS-G [Pan et al., 2024]: KOSMOS-G is a multimodal large language model designed for zero-shot image generation from interleaved vision-language inputs, including multiple images and text. It aligns the output space of a transformer-based causal language model with a diffusion-based image decoder using a lightweight AlignerNet and compositional instruction tuning. This architecture enables KOSMOS-G to perceive complex multimodal prompts and generate coherent, subject-driven images without modifying the base image decoder.
@@
  • Emu2 [Sun et al., 2024c]: Emu2 is a 37-billion-parameter generative multimodal model trained on large-scale multimodal sequences with a unified autoregressive objective. It exhibits strong in-context learning abilities for various multimodal tasks, including visual prompting and object-grounded generation.
@@
  • IP-Adapter [Ye et al., 2023]: IP-Adapter is a lightweight adapter that enables image prompt capability for pre-trained text-to-image diffusion models. It integrates image features into the generation process without modifying the base model, supporting flexible and efficient image-to-image generation.
@@
  • DreamEngine [Chen et al., 2025]: DreamEngine is a unified framework that integrates multimodal encoders with diffusion models through a two-stage training approach, enabling advanced text-image interleaved control and achieving state-of-the-art performance in generating images with complex, concept-merged inputs.
@@
  • Unified-IO 2 [Lu et al., 2023]: Unified-IO 2 is an autoregressive multimodal model capable of understanding and generating images, text, audio, and actions. It tokenizes various modalities into a shared semantic space and processes them with a single encoder-decoder transformer. Trained from scratch on a large multimodal pre-training corpus and fine-tuned on an ensemble of 120 datasets, Unified-IO 2 achieves state-of-the-art performance on the GRIT benchmark and strong results across more than 35 benchmarks.
@@
  • Lumina-mGPT [Zhuo et al., 2024]: Lumina-mGPT is a multimodal autoregressive models designed for flexible photorealistic text-to-image generation. It employs a pretrained decoder-only transformer as a unified framework for modeling multimodal token sequences. Through multimodal Generative PreTraining (mGPT) and subsequent Flexible Progressive Supervised Finetuning (FP-SFT) and Omnipotent Supervised Finetuning (Omni-SFT), Lumina-mGPT demonstrates versatile multimodal capabilities, including visual generation tasks, controllable generation tasks and vision-language tasks.
@@ F.1 Image Reconstruction
As illustrated in Figures 8 and 9, Mentor demonstrates strong image reconstruction capabilities following two-stage training. Notably, it is able to effectively reconstruct input images and preserve fine-grained visual details, even when input images are of low resolution (224×224) and outputs are generated at 512×512 resolution.
@@ F.3 Multi-Image Generation
We evaluate Mentor on multi-image generation tasks using the X2I dataset [Xiao et al., 2024]. As shown in Figure 11, the model is capable of generating visually consistent outputs conditioned the multi-image inputs. The generated images reflect coherent semantics, style, and layout across the samples.
@@ F.4 Multimodal In-Context Image Generation
To assess Mentor’s few-shot generalization capabilities, we evaluate it on the multimodal in-context image generation task using the X2I-ICL dataset [Xiao et al., 2024]. As illustrated in Figure 12, Mentor learns to synthesize images that follow the stylistic patterns demonstrated in the in-context examples. This indicates its capability to infer complex visual trends and align generation with image context.
@@ Appendix G Versatility Across Different Multimodal Tasks
We fine-tune the Stage 2 model on a subset of X2I-subject-driven [Xiao et al., 2024] dataset for two additional epochs using a reduced learning rate of $5\times 10^{-5}$. All other optimization settings remain consistent with Stage 2. The dataset is split into disjoint training and test sets, and quantitative results are reported on the test split. The model learns to generate visually diverse yet semantically aligned images for the same input.
@@
We fine-tune the model for 10 epochs on the X2I-ICL dataset [Xiao et al., 2024], which features sequences of input-output pairs for in-context generalization. We use a learning rate of $5\times 10^{-5}$ and ensure a strict train-test separation. The model adapts to context examples and generates new samples following the observed patterns, showing strong in-context learning performance without explicit prompting engineering.
@@ Conclusion.
Appendix H Limitations
Our work presents a promising alternative to diffusion-based methods for multimodal-conditioned image generation. As such, our focus is on evaluating performance under comparable conditions—i.e., with similar model capacities and training paradigms. However, the effectiveness of our approach is currently constrained by the limitations of available autoregressive (AR) backbone models. Due to the lack of high-performance AR generators, Mentor exhibits shortcomings in several aspects of image generation, including spatial reasoning, object counting, fine-grained human rendering, and stylization. These limitations reflect the current gap between current SOTA diffusion and autoregressive architectures in terms of generation fidelity and domain generalization. Additionally, while our training data is sourced from publicly available datasets and our synthetic data pipeline includes NSFW safeguards, a comprehensive evaluation of safety, fairness, and potential misuse remains lacking. Future work should incorporate thorough assessments of model biases and unintended behaviors. Finally, while our framework demonstrates strong versatility across diverse multimodal tasks, achieving competitive performance in specific domains may require more specialized training and the integration of more powerful multimodal encoders and generators. These initial findings nonetheless highlight the framework’s potential as a unified and efficient foundation for conditional multimodal image generation.

### v2

@@ E.1 Benchmark Details
Data Organization. DreamBench++ (Peng et al., 2025) comprises 150 high-quality reference images, sourced from Unsplash, Rawpixel, and Google Images, encompassing a balanced mix of subjects. These are evenly divided into three broad categories: objects, living subjects (humans and animals), and styles (illustrative, painterly, etc.), ensuring visual and conceptual diversity.
[…]
@@
The original DreamBench (Ruiz et al., 2023) dataset consists of 30 subjects, each paired with 25 prompts, totaling 750 prompt-image pairs. It serves as a foundational benchmark for evaluating personalized image generation models, focusing on the model’s ability to maintain subject identity across diverse prompts.
@@ E.2 Baselines
  • Textual Inversion (Gal et al., 2022) learns a new word embedding to represent a specific concept, enabling personalized image generation by incorporating the new token into prompts. It requires a few images of the subject and fine-tunes the embedding without altering the base model weights.
@@
  • DreamBooth (Ruiz et al., 2023): DreamBooth fine-tunes a pre-trained text-to-image model to bind a unique identifier with the subject’s visual concept, allowing for personalized generation. It requires several images of subject and modifies model weights to capture subject-specific details.
@@
  • BLIP-Diffusion (Li et al., 2023a): This approach introduces a pre-trained multimodal encoder to provide subject representations for the diffusion generator, enabling controllable multimodal image generation.
@@
  • KOSMOS-G (Pan et al., 2024): KOSMOS-G is a multimodal large language model designed for zero-shot image generation from interleaved vision-language inputs, including multiple images and text. It aligns the output space of a transformer-based causal language model with a diffusion-based image decoder using a lightweight AlignerNet and compositional instruction tuning. This architecture enables KOSMOS-G to perceive complex multimodal prompts and generate coherent, subject-driven images without modifying the base image decoder.
@@
  • Emu2 (Sun et al., 2024c): Emu2 is a 37-billion-parameter generative multimodal model trained on large-scale multimodal sequences with a unified autoregressive objective. It exhibits strong in-context learning abilities for various multimodal tasks, including visual prompting and object-grounded generation.
@@
  • IP-Adapter (Ye et al., 2023): IP-Adapter is a lightweight adapter that enables image prompt capability for pre-trained text-to-image diffusion models. It integrates image features into the generation process without modifying the base model, supporting flexible and efficient image-to-image generation.
@@
  • DreamEngine (Chen et al., 2025b): DreamEngine is a unified framework that integrates multimodal encoders with diffusion models through a two-stage training approach, enabling advanced text-image interleaved control and achieving state-of-the-art performance in generating images with complex, concept-merged inputs.
@@
  • Unified-IO 2 (Lu et al., 2023): Unified-IO 2 is an autoregressive multimodal model capable of understanding and generating images, text, audio, and actions. It tokenizes various modalities into a shared semantic space and processes them with a single encoder-decoder transformer. Trained from scratch on a large multimodal pre-training corpus and fine-tuned on an ensemble of 120 datasets, Unified-IO 2 achieves state-of-the-art performance on the GRIT benchmark and strong results across more than 35 benchmarks.
@@
  • Lumina-mGPT (Zhuo et al., 2024): Lumina-mGPT is a multimodal autoregressive model designed for flexible photorealistic text-to-image generation. It employs a pretrained decoder-only transformer as a unified framework for modeling multimodal token sequences. Through multimodal Generative PreTraining (mGPT) and subsequent Flexible Progressive Supervised Finetuning (FP-SFT) and Omnipotent Supervised Finetuning (Omni-SFT), Lumina-mGPT demonstrates versatile multimodal capabilities, including visual generation tasks, controllable generation tasks and vision-language tasks.
@@ F.1 Image Reconstruction
As illustrated in Figures 9 and 10, Mentor demonstrates strong image reconstruction capabilities following two-stage training. Notably, it is able to effectively reconstruct input images and preserve fine-grained visual details, even when input images are of low resolution (224×224) and outputs are generated at 512×512 resolution.
@@ F.2 Text-guided Image Segmentation
Open-Vocabulary Segmentation.
To further assess segmentation capability, we evaluate Mentor on ADE20K (A-150), a standard open-vocabulary segmentation benchmark. Notably, we perform zero-shot evaluation directly after Stage-1 pretraining, without any task-specific fine-tuning on these datasets. Due to the fixed output resolution of our generator, all input images are resized to $\times512$ for inference.
Figure 12: Qualitative results of open-vocabulary segmentation using Mentor on ADE20K (Zhou et al., 2017), PASCAL Context (Mottaghi et al., 2014), and PASCAL VOC (Everingham and Winn, 2011) benchmarks. The generated segmentation masks are visualized in red for clarity.
Figure 13: Qualitative examples from Image Editing. The model is finetuned on 1M samples from ImgEdit (Ye et al., 2025) and OmniEdit (Wei et al., 2024) dataset. Although restricted by the performance of image generator, Mentor can generate visually consistent edits.
Table 8: Zero-shot open-vocabulary segmentation on ADE20K-150 (Zhou et al., 2017). We report the mIoU score as the metric.
Dataset LSeg+ OpenSeg SimBaseline ZegFormer SimSeg DreamLIP Mentor
(Li et al., 2022) (Li et al., 2022) (Li et al., 2022) (Li et al., 2022) (Li et al., 2022) (Li et al., 2022)
ADE20K-150 13.00 15.30 15.30 16.40 20.50 17.10 19.91
As shown in Table 8 and Figure 12, Mentor achieves competitive performance on the benchmark compared with the traditional open-vocabulary segmentation methods in a zero-shot manner. Compared with these methods that require massive task-specific pretraining, Mentor shows competitive performance after Stage-1 training in a zero-shot manner. These results demonstrate that our two-stage training effectively learns to balance visual and textual guidance, enabling the model to generate accurate segmentation masks for novel categories. While performance is constrained by the current backbone capacity and data scale, these findings highlight the extensibility and promising potential of our framework for broader vision-language tasks.
@@ F.3 Multi-Image Generation
We evaluate Mentor on multi-image generation tasks using the X2I dataset (Xiao et al., 2024b). As shown in Figure 14, the model is capable of generating visually consistent outputs conditioned on the multi-image inputs. The generated images reflect coherent semantics, style, and layout across the samples.
@@ F.4 Multimodal In-Context Image Generation
To assess Mentor’s few-shot generalization capabilities, we evaluate it on the multimodal in-context image generation task using the X2I-ICL dataset (Xiao et al., 2024b). As illustrated in Figure 15, Mentor learns to synthesize images that follow the stylistic patterns demonstrated in the in-context examples. This indicates its capability to infer complex visual trends and align generation with image context.
@@ Appendix G Versatility Across Different Multimodal Tasks
We fine-tune the Stage 2 model on a subset of X2I-subject-driven (Xiao et al., 2024b) dataset for two additional epochs using a reduced learning rate of $5\times 10^{-5}$. All other optimization settings remain consistent with Stage 2. The dataset is split into disjoint training and test sets, and quantitative results are reported on the test split. The model learns to generate visually diverse yet semantically aligned images for the same input.
@@
We fine-tune the model for 10 epochs on the X2I-ICL dataset (Xiao et al., 2024b), which features sequences of input-output pairs for in-context generalization. We use a learning rate of $5\times 10^{-5}$ and ensure a strict train-test separation. The model adapts to context examples and generates new samples following the observed patterns, showing strong in-context learning performance without explicit prompt engineering.
@@ Conclusion.
Appendix H Human Evaluation
To validate the reliability of GPT-based evaluation metrics used in DreamBench++, we conduct a human study on 50 randomly sampled images from our model’s outputs. Three independent annotators rate each image on Concept Preservation (CP) and Prompt Following (PF) using the same five-level ordinal scale (0–4) as the automatic evaluation.
We report two complementary measures of reliability:
  • Inter-annotator agreement: Measured by Krippendorff’s $\alpha$, we obtain $\alpha=0.76$ for CP (substantial agreement) and $\alpha=0.62$ for PF (moderate agreement). These values indicate consistent human judgment across annotators.
  • Human-metric consistency: Defined as the proportion of cases where the GPT-based score matches the majority human judgment, this metric reaches 84% for CP and 94% for PF. These figures align closely with the consistency rates reported in DreamBench++ (83.31% for CP, 98.17% for PF).
These results confirm that the GPT-based metrics serve as reliable proxies for human evaluation, supporting their use as the primary evaluation protocol in our experiments.
Appendix I CP–PF Trade-off Analysis
A central challenge in multimodal-conditioned image generation is balancing fidelity to the reference image (Concept Preservation, CP) against adherence to the text prompt (Prompt Following, PF). In this section, we conduct systematic experiments to analyze the CP–PF trade-off by varying two key factors: (1) Stage-2 task mixing ratios during training, and (2) classifier-free guidance (CFG) scale at inference.
I.1 Effect of Task Weight Distribution
We investigate how the composition of training tasks in Stage-2 affects the balance between concept preservation and prompt following. Stage-2 comprises four task types with configurable sampling weights: text-to-image generation (T2I), object segmentation (Seg), image recovery (Rec), and subject-driven generation (Sub). We evaluate five representative weight configurations, as reported in Table 9.
Table 9: Effect of Stage-2 task weight distribution on DreamBench++. Left columns show sampling weights for each task type; right columns report evaluation metrics.
Task Sampling Weights Metrics
[…]
Default (Ours) 0.45 0.15 0.15 0.25 0.557 0.840 0.468
Image-heavy 0.20 0.25 0.25 0.30 0.609 0.741 0.451
Text-heavy 0.80 0.05 0.05 0.10 0.383 0.915 0.350
w/o Rec 0.45 0.25 0.05 0.25 0.588 0.666 0.392
w/o Seg 0.45 0.05 0.25 0.25 0.390 0.880 0.343
The results reveal a systematic trade-off between concept preservation and prompt following:
  • Image-heavy: Increasing weights for image-centric tasks (Seg, Rec, Sub) improves CP from 0.557 to 0.609 but reduces PF from 0.840 to 0.741.
  • Text-heavy: Emphasizing T2I (weight = 0.80) yields higher PF (0.915) at the cost of significantly degraded CP (0.383).
  • w/o Rec: Reducing image recovery weight primarily degrades PF and induces copy-paste behavior.
  • w/o Seg: Minimizing segmentation weight mainly hurts CP, impairing fine-grained visual grounding.
[…]
I.2 Effect of Classifier-Free Guidance Scale
We examine how classifier-free guidance (CFG) scale at inference time affects the CP–PF trade-off. Table 10 summarizes the results across four CFG values.
Table 10: Effect of CFG scale on DreamBench++ metrics. Higher CFG strengthens prompt adherence but may degrade concept preservation at extreme values.
CFG Scale 3.5 5.5 7.5 9.5
[…]